OpenStudy (anonymous):

What is the slope of the line 4x – 8y – 16 = 0 in simplest form? ?/?

OpenStudy (anonymous):

in xyz form right?

OpenStudy (xapproachesinfinity):

first rewrite that in slope intercept form \[y=mx+b\] you can do this by solving for y

OpenStudy (anonymous):

fan and medal.

OpenStudy (xapproachesinfinity):

or just divide everything by -8 tp get \[\frac{-1}{2}x+y+2=0\] so the slope is -1/2

OpenStudy (xapproachesinfinity):

i mean 1/2

OpenStudy (xapproachesinfinity):

do you get it?

OpenStudy (xapproachesinfinity):

if you did the other way solving for y you should have \[y=\frac{1}{2}x-2\] so m=1/2 m is the slope

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

so then the answer would be -1/2